OpenStudy (gabylovesyou):

Convert the equation you arrived at in question 2 into slope-intercept form. Make sure to include all of your work. equation i did in question 2 was " x + y = 3 " @jim_thompson5910

OpenStudy (dtan5457):

put x + y = 3 into slope intercept?

OpenStudy (gabylovesyou):

yes

OpenStudy (dtan5457):

Slope, point = y-y1=m(x-x1) the point you know for sure is here is 3,0 because that's the point of the y intercept. x+y=3 y=3-x slope=-x Plug all that info into the formula

OpenStudy (gabylovesyou):

im confused

OpenStudy (dtan5457):

y-y1=m(x-x1) m=slope x1,y1=3,0

OpenStudy (dtan5457):

Just plug it in

OpenStudy (gabylovesyou):

how do u know its 3,0

OpenStudy (dtan5457):

y intercept=3 3,0

OpenStudy (dtan5457):

as one of the line's points

OpenStudy (gabylovesyou):

so then its y - 0 = m(x - 3)

OpenStudy (dtan5457):

m=-1

OpenStudy (dtan5457):

y-0=-1(x-3)

OpenStudy (gabylovesyou):

how do u know that ? o-o

OpenStudy (dtan5457):

x + y = 3 that's the same thing as y=3-x which is in slope intercept form. -x=-1

Nnesha (nnesha):

" x + y = 3 " this one into slope intercept form ??

OpenStudy (gabylovesyou):

@dtan5457 ohhh @Nnesha yes

Nnesha (nnesha):

just one step to convert this into slope intercept form

Nnesha (nnesha):

slope intercept form is y = mx +b where m is slope which is always with x and y is y-intercept

Nnesha (nnesha):

so you have to solve for y x+y = 3 solve for y

OpenStudy (dtan5457):

lol wow. I put it into slope point for no reason, I thought that's what the question asked...I need to read better.

Nnesha (nnesha):

;)

Nnesha (nnesha):

so can you solve for y x+y = 3 ???

OpenStudy (gabylovesyou):

how...

OpenStudy (dtan5457):

subtract the x.

Nnesha (nnesha):

x is positive on left side you have to move x on right side so do opposite of postivie subtract x both side

OpenStudy (gabylovesyou):

y = 3 - x ?

Nnesha (nnesha):

perfect

Nnesha (nnesha):

now you can wriet like this y = -x+3 x first so its looks good

OpenStudy (gabylovesyou):

lol ok

OpenStudy (gabylovesyou):

thats it ?

OpenStudy (dtan5457):

yep

OpenStudy (gabylovesyou):

oh psh. lol cx THANK YOU

OpenStudy (dtan5457):

slope intercept is y=mx+b y=-x+3 is in that form

Nnesha (nnesha):

\(\Huge \color{orange}{\star^{ \star^{\star:)}}}\Huge \color{purple}{\star^{ \star^{\star:)}}}\) \(\huge\color{green}{{\rm my }~\rm pleasure !! :)}\) \(\Huge \color{orange}{\star^{ \star^{\star:)}}}\Huge \color{purple}{\star^{ \star^{\star:)}}}\)

Nnesha (nnesha):

nvm jk